OpenStudy (anonymous):

please help :( can someone please solve and graph |2x-5|+3 less than or equal to 10

geerky42 (geerky42):

You need to solve it by graphing?

OpenStudy (anonymous):

Nopee i need to solve then graph

geerky42 (geerky42):

Okay, Best thing to do first is to isolate absolute value, right? Just subtract both sides by 3, then you should have \(|2x-5|\le7\). Any idea on how to get rid of absolute value?

OpenStudy (anonymous):

i think i add and subtract 5?

geerky42 (geerky42):

Well, that won't help get rid of it, though. \[|2x-5|+5\le12\]

OpenStudy (anonymous):

i thought i leave the equation as it is after subtracting 3. and then solve it but also do another equation with 2x+5

geerky42 (geerky42):

Yeah kind of. Let just say \(2x+5 =a\) So here you have \(|a|\le7\). Do you know what value \(a\) only can be?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

?

geerky42 (geerky42):

Well, yeah \(7\le7\) But there are more, right? what about 6? 5? and so on?

geerky42 (geerky42):

Until 0?

geerky42 (geerky42):

Can we go behind 0 to negative number?

geerky42 (geerky42):

How far can we go?

OpenStudy (anonymous):

oh ok well i really just need help on this specific problem being that i have many more to do due by 10

geerky42 (geerky42):

Well, I'm just trying to help you understand how to solve it, so you don't have to ask similar questions again and again.

OpenStudy (anonymous):

of course thanks

geerky42 (geerky42):

But if you need answer for grade sake, well, sorry to say but we aren't allowed to do that lol...

geerky42 (geerky42):

Can you spare time?

OpenStudy (anonymous):

that isnt what i was asking for

geerky42 (geerky42):

Okay, back to stuffs. \(|a|\le7\) What range of number can \(a\) be?

geerky42 (geerky42):

It only can goes from 7 to -7, right?

geerky42 (geerky42):

Because \(|-7.1|\le7\) isn't correct, right?

geerky42 (geerky42):

So you can "translate" \(|a|\le 7\) to \(-7\le a\le7\) Makes sense so far?

OpenStudy (anonymous):

mhmm

geerky42 (geerky42):

So here \(|2x-5|\le 7\quad\Longrightarrow\quad -7\le2x-5\le 7\)

geerky42 (geerky42):

Now you can do algebra from here

OpenStudy (anonymous):

idk how to graph x\[x \le6,1\]

geerky42 (geerky42):

How did you get to \(x\le6,1\)?

OpenStudy (anonymous):

i solved it dude

OpenStudy (anonymous):

im sorry i dont have time for a lesson at the moment so im trying to rush and use prior knowledge

geerky42 (geerky42):

It's \(-1\le x\le 6\) This mean x can only be in between -1 to 6 inclusive. So you can graph like this: